# https://cs.uwaterloo.ca/~smwatt

### v1

@@ 1 Introduction
Software systems for computer algebra have been available since the 1960s on a wide variety of computing platforms. The underlying architectures of these platforms varied in major respects and numerous challenges had to be overcome to allow a given computer algebra system to operate in different environments. Computer algebra programs have characteristics that place particular demands on the supporting software: dynamic memory allocation was required to handle a great many small, mostly read-only, data objects with varying lifetimes and the efficiency of certain key data types (such as arbitrary precision integers or integers mod a prime) was absolutely critical to overall system performance. For this reason, the software portability issues for computer algebra have a different emphasis than is found elsewhere. This article lays out the issues we have observed in the creation of computer algebra systems, and the open issues we see remaining.
@@ 2 The Meaning of Portability
In the early days of computer algebra (notably much of the 1960s) serious scale algebraic computation strained (at the very least!) machine capacity, and porting only made sense if the new target hardware was of a style installed at multiple or particularly important sites. A simplistic characterization was that only the larger machines from IBM, DEC and Cray mattered. But the designs from those three major vendors had little in common either in hardware architecture or standard software. Programming language standardization emerged through the 1960s with FORTRAN[19] leading the way and COBOL following (though probably little serious computer algebra has used COBOL), but even beyond the release of the standard it would both be necessary to make careful checks about the compliance level of a local compiler and to be quite thoughtful about the extent to which tempting local extensions would make code run faster or better. The other viable path was to use Lisp where although there was good agreement about the specification of its core, versions at different sites or on different computers were enthusiastically enhancing and adjusting their version for local needs. The section we have later on using Reduce as a case-study may give more insight into those past and dark times.
@@
A number of years later Unix had become a dominant operating system for use in the sort of research environment where algebra systems were being developed. C was ubiquitous and Lisp had gained a dominant position for symbolic computation including but going well beyond algebra. However there had been an explosion in the number of vendors of workgroup-scale computers and neither Unix nor C nor Lisp had internationally accepted standard that could be relied on. Anybody who lived through those decades is likely to recall that making C code portable involved scattering horrendous numbers of “#ifdef” directives through the code. Some would be to allow for operating environment issues, others for word-length, byte order, floating point representation and other hardware matters, or because one idiom was much faster on platform A than B while the situation was reversed for another code fragment that was seen as performance critical. Finally, it was not uncommon to need to use conditional compilation to sidestep bugs in the compiler or some of the libraries that were to be used. Maple[36] is an example of a system that emerged against this background. Its history illustrates the challenge of the cost of the computers used elsewhere to develop computer algebra, leading the Maple designers keen to be able to exploit smaller and more affordable options. It also shows how they based what they were doing on C (with very initial work using its predecessor languages). And despite the fact that C had an ad hoc rather than a formal standard and no doubt with a serious amount of conditional compilation they made their system so it could be transplanted to a range of Unix platforms: portability was in their minds from the start.
@@
Towards the end of the 1980s a standard for C emerged and POSIX[16] at least explained what a Unix system ought to support. Reworking code to be in ANSI C and so that it relied on POSIX promised to greatly ease the task of migrating code between machines from different vendors — and this just in time to see the start of a need to cope with the emerging personal computers that had become powerful enough that and individual scientist could often achieve as much using one on that was their own and could be left running overnight if necessary as they could on a more powerful departmental or institutional machine. In our discussion of Aldor[49] and the Axiom[27] system it was associated with, we note that at the time NAG Ltd took over responsibility for marketing Axiom it could run on personal-scale hardware at least in cases where the machine was generously (for the time) endowed with memory. So as variants on Linux really started to converge making life easier for developers and maintenance teams, the joys of Widows and the Macintosh emerged and the feeling that users now needed comfortable graphical interfaces (where portable support was even more “exciting”) arose.
@@
The last remark above highlights a separate issue that can be held to fall under the banner of portability. Early system of necessity used text-based interfaces [^1] Apart from some really good early work at MIT and Xerox PARC where special hardware was available to the few.. When those systems moved into a modern world it was necessary to enhance them with GUIs. While from some perspective one can views this as an extension rather than as porting, our view that it falls into the same challenge and the workload will tend to fall on the same sort of people would would migrate the code to a new computer. It is porting across the boundaries of time and user expectation rather than across boundaries of operating system dependencies. Other rather similar challenges arose with a need for web-based interfaces and it is plausible that two pressing ones today are to find a response to generative AI and to provide full exploitation of the concurrency opportunities of cloud-based servers. Each of these can call for re-work deep into the core of a project.
@@
Just when you had hoped that portability challenges were receding into the past, with the transition from 32-bit wordlength to 64-bits having been one of the more notable (that caused some entertainment for those Reduce developers who were concerned with the lowest level of the software stack), one finds two disruptions. The first is when major commonly used brands of computer change the style of CPU they use. Supporting the new situation will of course be merely a variant on the traditional “port to a different system” activity, so those who had been active in the 1980s will cope well. However experience shows that support libraries and the like for mildly fringe work such as that on algebra systems can take a significant time to stabilise. Also re-working build scripts and the schemes for packing software for distribution such that recipients who have both old and new architectures both have a seamless experience takes effort and sometimes really detailed work. Of course the computer vendors in presenting their transition tend to promise that the adjustments will be really easy to make: experience qualifies this stance — it is liable to spiritually easy but getting everything sorted out can take an astonishing amount of time.
@@
A quite different style of porting is called for when some significant intellectual development has been made and the originator created an implementation of it. They will typically have done that as an extension to or as code in the context of some existing algebra system. They might well make their code available to all under some sort of generous and flexible license. But those who would like to either test and explore the work or leverage it to enhance the capabilities of their own system may find they have something of a porting mountain to climb. We can cite two separate cases of this that we are interested in which both happen to involve indefinite integration. They also both happen to be cases where the originator of the code of interest is deceased and hence not in a position to assist:
@@
    Manual Bronstein worked on systematic schemes for integration[3], building on the work on Risch[43]. His implementation for Axiom (see our section on Aldor) is probably the most complete development of the mathematical basis for the task. However now not only is Bronstein not with us, but the major developer who has been working on that for many years appears to have given up. So there is a serious issue of portability that has to be overcome to arrange that the practical aspects of Bronstein’s work being lost forever. An optimistic response to this is to consider AI transformation from the code expressed in Axiom’s rather particular language into something more widely used;
@@
    Albert Rich had been taking an almost diametrically opposite approach to indefinite integration, and had been developing a very extensive rule-based system called RUBI[26]. Part of the desire there was not just to evaluate the integrals but to generate results in concise and tidy form — something that schemes based on the formal mathematical approach often fail to achieve. He had coded his rule-set using Mathematica[18], and using it there it performs remarkably well on various test suites. He made his rules freely available, but when one attempts to use them free-standing it gradually becomes clear that at least in the form that he left them in there is much reliance not just on Mathematica syntax (which is rather easy to reproduce) but on the detailed semantics of the pattern matching there and the incidental simplification that RUBI expects Mathematica to do for it. The code there is still available and can be useful for those who have a Mathematica license (or for that company to embed within their commercial system), but porting it out for use elsewhere has proved non-trivial[40].
When significantly mathematical code is to be ported then there are a wide range of ways in which different systems interpret similar notations in different ways. This can include the treatment of multi-valued functions and indeed whether formulae are to be viewed as existing in the real of complex domain or some more complicated world. In terms of communication with users $x\sqrt{x}$ and $x^{3/2}$ may alter the appearance of formulae enough to be practically important even if mathematically fairly trivial. Use of mathematical notation is notably flexible and in different contexts conflicting conventions can apply. As a simple concrete instance, the Bernoulli number with index 1 may be either $+1/2$ or $-1/2$ depending on the use to be made of it! We elaborate one some more of these challenges in Section 8.
@@
The very essence of the scientific method is that when an advance has been made others should be able to reproduce and validate it. In computer algebra’s early days the papers presenting new algorithms showed techniques that the reader could fairly readily re-implement, but even then a number of them included the code that the authors had produced. These days, when code is usually way to complicated to recreate from a mere description, it is commonplace for code to be deposited on some publicly available archive. If that code will only work using exactly the computer configuration and support software context that the author enjoyed then even in the short term its usefulness to others may be constrained, and very typically over time as that setup becomes out of date it will lose value. In the case of a package to perform some algebraic transformation in a way that moves the state of the art forward there can be a level of ambiguity as between the new work and the strengths of the platform it is built on, so other workers could reasonably want to try it atop a different base. The issues of being able to do this alongside the ones forced by continuing changes in the world of computers puts portability at the heart of all we do. The authors here have perhaps been particularly been brought face to face with this issue by the examples regarding indefinite integration cited later on where in one case their own earlier contribution had included integration code that needs to be updated or replaced and in the other where there was direct involvement in the creation of the programming system used in one of the projects considered – but where the whole language now looks at risk.
Sometimes making software portable adds a non-trivial burden to the implementer, whether it be a 1960’s FORTRAN programming needing to know about “definition at the second level[35]” or today’s C++ users about “strict aliasing[53]” and many other easy to use idioms that can lead to unexpected behaviour on some but not all platforms. There can also be coding patterns that may lead to widely varying system-dependent performance (verilog can exhibit this in spades, but those who have worked in high performance computing will also be very aware that tuning can lead to dramatic performance shifts). However for many people whose career path depends mainly on their publication record the incentive can be to get an implementation of their ideas working well enough on just the computer they happen to be using this month, report on results and timings and move on — leaving their code available to all to download but without true certainty that it will remain usable either elsewhere or in the future.
@@
    Those building an artifact that they hope very many people will use and find so central to their process that they cite its originator, leading to the distribution of the software being viewed as having high impact. Both the early work on Reduce and Maple took this stance;
@@
    Employees in a company preparing commercial software where portability across existing platforms and to potential future ones may broaden the market. Maple and Mathematica must of necessity followed this pattern;
@@
We fall into the first category when we find that against plausibility software that we developed many years ago is still in use, and we look back in horror at some of the coding styles we originally used and but feel that keeping it current extends the pleasure we get from knowing that our code is used to solve real problems. But more than that we are in the third category where we take pride in keeping the software we support going on older systems nut also arranging it can take advantage of whatever the latest platform it. In that context we also believe that if open source software is to attract a new generation of maintainers the code should look tidy (and ideally this associates with portability) and is not limited by only running on archaic platforms.
@@ 3 Early Portability in Computer Algebra
Alpak was rewritten as a set of Fortran subroutines and paired with a language translator to become Altran[31] whose portability is mentioned by Johnson and Richie[30] when they were considering how best to make C and Unix easier to move from system to system.
@@
The Camal[11] computer algebra system, developed by David Barton, Steve Bourne and John Fitch, was originally written in assembly code for the Titan computer at Cambridge. There were just two very similar computers in the world, and another three that were fairly close relatives..
Over the same period, George E. Collins was developing a set of libraries[5] for big integer and polynomial arithmetic, first at IBM, then at the University of Wisconsin. We are under the impression that these began as assembly code libraries before the development shifted to Fortran, but would appreciate confirmation of this.
@@
Two alternative approaches emerged. The first was to rely on Lisp[32] as an implementation language, as Lisp had the concept of dynamic data structures at its core. The second was to rely on lower-level “systems programming languages” to provide efficient resource management. At the start this meant BCPL[42], although in due course C[33] claimed the dominant place in that ecological niche.
From its beginnings at the end of the 1950s, Lisp had matured and taken its place as a recognised family of closely related languages. There was no notion of a single, standard Lisp language. Important variants around 1970 were MacLisp[41] (from MIT) and InterLisp[46] (then from Bolt, Beranek and Newman). For some time, both of those had a clear enough common core that migrating software between them was reasonably sane for quite a while.
However by the end of the 1960s dialect divergence was sufficient to cause trouble and, Anthony Hearn introduced Standard Lisp[15] to serve as the foundation for the Reduce system. This was not so much intended for adoption outside Reduce usage, but it defined the expectations that Reduce had. It was consciously conservative and a bit minimal, and hence it was not especially attractive to those who wanted to develop and extent Lisp for direct use. So meanwhile the Macsyma system was built on MacLisp, first led by William Martin and then Joel Moses from 1968 to 1982 at MIT, and the Lisp dialect used there collected many many addons, where Joel Moses characterised the language as “a ball of mud”[38] that one could add arbitrary extras to without altering its already slightly odd outline. Later, a port of Macsyma to the VAX architecture was achieved at UC Berkeley by creating a MacLisp-like base, Franz Lisp.
The Scratchpad system was built on Lisp/VM[2], led by Jim Griesmer at IBM. The IBM Lisp system was led by Fred Blair and later Cyril Alberga. (Note, this Scratchpad was retrospectively named Scratchpad I to distinguish it from the later, distinct system Scratchpad II which became Axiom.). Scratchpad (I) was seriously notable from the portability perspective because it started life by amalgamating major components from almost all of the Lisp-coded algebra projects extent in around 1970 when the project was in its early stages the various Lisp dialects were still close enough that code for one could be adapted for use on another with only a reasonable amount of effort.
@@
The different Lisp systems were not equally available on all platforms, so some systems made an attempt to run on more than one Lisp, despite the differences. Reduce is particularly notable in this regard. A version of it dated 1973 has been recovered and it can still be run — and amazingly much of the code in it can be recognised in the most recent versions of the system. This historical relic is available for download on the main Reduce website at https://sourceforge.net/projects/reduce-algebra. The strategy there was to rely on a minimal set of core Lisp functionality, and to reproduce necessary higher-level function as as a set of Lisp functions relying only on that core. A similar approach was later adopted for Axiom.

### v4

# https://cs.uwaterloo.ca/˜smwatt

@@ 1 Introduction
This article lays out the issues we have observed in the creation of computer algebra systems, and the open issues we see remaining. Software systems for computer algebra have been available since the 1960s on a wide variety of computing platforms. The underlying architectures of these platforms varied in major respects and numerous challenges had to be overcome to allow a given computer algebra system to operate in different environments. Computer algebra programs have characteristics that place particular demands on the supporting software: dynamic memory allocation was required to handle a great many small, mostly read-only, data objects with varying lifetimes and the efficiency of certain key data types (such as arbitrary precision integers or integers mod a prime) was absolutely critical to overall system performance. Another feature of computer algebra systems is their embodiment of multiple sophisticated algorithms, each involving numerous special cases (for efficiency) and implemented sitting atop a library of equally elaborate sub-algorithms. These algorithms cannot be readily moved from one system to another, so systems tend to be long-lived, spanning multiple generations of computing environments. For these reasons, the software portability issues for computer algebra have a different emphasis than is found elsewhere.
@@ 2 The Meaning of Portability
In the early days of computer algebra (notably much of the 1960s) serious scale algebraic computation strained (at the very least!) machine memory capacity, and porting only made sense if the new target hardware was of a style installed at multiple or particularly important sites. A simplistic characterisation was that only the larger machines from IBM, DEC and Cray mattered. But the designs from those three major vendors had little in common either in hardware architecture or standard software. Programming language standardisation emerged through the 1960s with Fortran[21] leading the way and Cobol following (though probably little serious computer algebra has used Cobol), but even beyond the release of the standard it would both be necessary to make careful checks about the compliance level of a local compiler and to be quite thoughtful about the extent to which tempting local extensions would make code run faster or better. The other viable path was to use Lisp where although there was good agreement about the specification of its core, versions at different sites or on different computers were enthusiastically enhancing and adjusting their version for local needs. The section we have later on using Reduce as a case-study may give more insight into those past and dark times.
@@
A number of years later Unix had become a dominant operating system for use in the sort of research environment where computer algebra systems were being developed. C was ubiquitous and Lisp had gained a dominant position for symbolic computation including but going well beyond algebra. However there had been an explosion in the number of vendors of workgroup-scale computers and neither Unix nor C nor Lisp had internationally accepted standards that could be relied on. Anybody who lived through those decades is likely to recall that making C code portable involved scattering horrendous numbers of “#ifdef” directives through the code. Some would be to allow for operating environment issues, others for word-length, byte order, floating point representation and other hardware matters, or because one idiom was much faster on platform A than B while the situation was reversed for another code fragment that was seen as performance critical. Finally, it was not uncommon to need to use conditional compilation to sidestep bugs in the compiler or some of the libraries that were to be used. Maple [37] is an example of a system that emerged against this background. Its history illustrates the challenge of the cost of the computers used elsewhere to develop computer algebra, leading the Maple designers to be keen to exploit smaller and more affordable options. It also shows how they based what they were doing on C (with very initial work using B, its predecessor language). And despite the fact that C had an ad hoc rather than a formal standard and no doubt with a serious amount of conditional compilation they made their system so it could be transplanted to a range of Unix platforms: portability was in their minds from the start – indeed, it was in the title of the first Maple paper [7].
@@
Towards the end of the 1980s a standard for C emerged and Posix[19] at least explained what a Unix system ought to support. Reworking code to be in Ansi C and so that it relied on Posix promised to greatly ease the task of migrating code between machines from different vendors — and this just in time to see the start of a need to cope with the emerging personal computers. These had become powerful enough that an individual scientist could often achieve as much using one that was their own, and could be left running overnight if necessary, as they could on a more powerful departmental or institutional machine. In our discussion of Aldor[52] and the Axiom [29] system it was associated with, we note that at the time NAG Ltd took over responsibility for marketing Axiom, it could run on personal-scale hardware at least in cases where the machine was generously (for the time) endowed with memory. So as variants on Linux really started to converge making life easier for developers and maintenance teams, the joys of Windows and the Macintosh emerged and the feeling that users now needed comfortable graphical interfaces (where portable support was even more “exciting”) arose.
@@
The last remark above highlights a separate issue that can be held to fall under the banner of portability. Early systems of necessity used text-based interfaces [^1] Apart from some really good early work at MIT and Xerox PARC where special hardware was available to the few.. When those systems moved into a modern world it was necessary to enhance them with GUIs. While from some perspective one can views this as an extension rather than as porting, our view that it falls into the same challenge and the workload will tend to fall on the same sort of people who would migrate the code to a new computer. It is porting across the boundaries of time and user expectation rather than across boundaries of operating system dependencies. Other rather similar challenges arose with a need for web-based interfaces and it is plausible that two pressing ones today are to find a response to generative AI and to provide full exploitation of the concurrency opportunities of cloud-based servers. Each of these can call for re-work deep into the core of a project.
@@
Just when you had hoped that portability challenges were receding into the past, with the transition from 32-bit word length to 64-bits having been one of the more notable (that caused some entertainment for those Reduce developers who were concerned with the lowest level of the software stack), one finds two disruptions. The first is when major commonly used brands of computer change the style of CPU they use (e.g. Apple changing from PowerPC to Intel to ARM). Supporting the new situation will of course be merely a variant on the traditional “port to a different system” activity, so those who had been active in the 1980s will cope well. However experience shows that support libraries and the like for mildly fringe work such as that on computer algebra systems can take a significant time to stabilise. Also re-working build scripts and the schemes for packing software for distribution such that recipients who have both old and new architectures both have a seamless experience takes effort and sometimes really detailed work. Of course the computer vendors in presenting their transition tend to promise that the adjustments will be really easy to make: experience qualifies this stance — it is liable to theoretically easy but getting everything sorted out can take an astonishing amount of time.
@@
A quite different style of porting is called for when some significant intellectual development has been made and the originator created an implementation of it. They will typically have done that as an extension to or as code in the context of some existing computer algebra system. They might well make their code available to all under some sort of generous and flexible license. But those who would like to either test and explore the work or leverage it to enhance the capabilities of their own system may find they have something of a porting mountain to climb. We can cite two separate cases of this that we are interested in which both happen to involve indefinite integration. They also both happen to be cases where the originator of the code of interest is deceased and hence not in a position to assist:
@@
    Manuel Bronstein worked on systematic schemes for symbolic integration (antiderivatives) [5], building on the work on Risch[44]. His implementation for Axiom (see our section on Aldor) is probably the most complete development of the mathematical basis for the task. However now not only is Bronstein not with us, but the major developer who has been working on that for many years appears to have given up. So there is a serious issue of portability that has to be overcome so that the practical aspects of Bronstein’s work being lost forever. An optimistic response to this is to consider AI transformation from the code expressed in Axiom’s rather particular language into something more widely used;
@@
    Albert Rich had been taking an almost diametrically opposite approach to indefinite integration, and had been developing a very extensive rule-based system called Rubi[28]. Part of the desire there was not just to evaluate the integrals but to generate results in concise and tidy form — something that schemes based on the formal mathematical approach often fail to achieve. For example, in its most recent version, https://12000.org shows that given the integral $\int\frac{dx}{x^{5}(a+bx^{8})}$ (test number 245), Mathematica [54] produces output
    (-Sqrt[a] + Sqrt[b]*x^4*ArcTan[Cot[Pi/8] - (b^(1/8)*x*Csc[Pi/8])/a^(1/8)]
    + Sqrt[b]*x^4*ArcTan[Cot[Pi/8] + (b^(1/8)*x*Csc[Pi/8])/a^(1/8)]
    + Sqrt[b]*x^4*ArcTan[(b^(1/8)*x*Sec[Pi/8])/a^(1/8) - Tan[Pi/8]]
    - Sqrt[b]*x^4*ArcTan[(b^(1/8)*x*Sec[Pi/8])/a^(1/8) + Tan[Pi/8]])
    /(4*a^(3/2)*x^4)
    while Rubi gives
    (-(1/(a*x^4)) - (Sqrt[b]*ArcTan[(Sqrt[b]*x^4)/Sqrt[a]])/a^(3/2))/4
    Rich had coded his rule-set using Mathematica, and using it in that system it performs remarkably well on various test suites. He made his rules freely available, but when one attempts to use them free-standing it gradually becomes clear that, at least in the form that he left them in, there is much reliance not just on Mathematica syntax (which is rather easy to reproduce) but on the detailed semantics of the pattern matching there and the incidental simplification that Rubi expects Mathematica to do for it. The code there is still available and can be useful for those who have a Mathematica license (or for that company to embed within their commercial system), but porting it out for use elsewhere has proved non-trivial[41].
When significantly mathematical code is to be ported then there are a wide range of ways in which different systems interpret similar notations in different ways. This can include the treatment of multi-valued functions and indeed whether formulae are to be viewed as existing in the real or complex domain or some more complicated world. In terms of communication with users $x\sqrt{x}$ and $x^{3/2}$ may alter the appearance of formulae enough to be practically important even if mathematically fairly trivial. Use of mathematical notation is notably flexible and in different contexts conflicting conventions can apply. As a simple concrete instance, the Bernoulli number with index 1 may be either +1/2+1/2 or 1/2-1/2 depending on the use to be made of it! We elaborate on some more of these challenges in Section 8.
@@
The very essence of the scientific method is that when an advance has been made others should be able to reproduce and validate it. In computer algebra’s early days the papers presenting new algorithms showed techniques that the reader could fairly readily re-implement, but even then a number of them included the code that the authors had produced. These days, when code is usually way too complicated to recreate from a mere description, it is commonplace for code to be deposited on some publicly available archive. If that code will only work using exactly the computer configuration and support software context that the author enjoyed then even in the short term its usefulness to others may be constrained, and very typically over time as that setup becomes out of date it will lose value. In the case of a package to perform some algebraic transformation in a way that moves the state of the art forward there can be a level of ambiguity as between the new work and the strengths of the platform it is built on, so other workers could reasonably want to try it atop a different base. The issues of being able to do this alongside the ones forced by continuing changes in the world of computers puts portability at the heart of all we do. The authors here have perhaps been particularly been brought face to face with this issue by the examples regarding indefinite integration cited later on, where in one case their own earlier contribution had included integration code that needs to be updated or replaced and in the other where there was direct involvement in the creation of the programming system used in one of the projects considered – but where the whole language now looks at risk.
Sometimes making software portable adds a non-trivial burden to the implementer, whether it be a 1960’s Fortran programmer needing to know about “definition at the second level”[36] or today’s C++ users about “strict aliasing” [57] and many other easy to use idioms that can lead to unexpected behaviour on some, but not all, platforms. There can also be coding patterns that may lead to widely varying system-dependent performance (Verilog can exhibit this in spades, but those who have worked in high performance computing will also be very aware that tuning can lead to dramatic performance shifts). However, for many people whose career paths depend mainly on their publication record, the incentive can be to get an implementation of their ideas working well enough on just the computer they happen to be using this month, report on results and timings and move on — leaving their code available to all to download but without true certainty that it will remain usable either elsewhere or in the future.
@@
    Those building an artefact that they hope many people will use and find so central to their process that they cite its originator, leading to the distribution of the software being viewed as having high impact. Both the early work on Reduce and Maple took this stance;
@@
    Employees in a company preparing commercial software where portability across existing platforms and to potential future ones may broaden the market. Maple and Mathematica must have, of necessity, followed this pattern;
@@
We fall into the first category when we find that, against plausibility, software that we developed many years ago is still in use, and we look back in horror at some of the coding styles we originally used and but feel that keeping it current extends the pleasure we get from knowing that our code is used to solve real problems. But more than that we are in the third category where we take pride in keeping the software we support going on older systems but also arranging it can take advantage of whatever the latest platform it. In that context we also believe that if open source software is to attract a new generation of maintainers, the code should look tidy (and ideally this associates with portability) and is not limited by only running on archaic platforms. Documentation can be helpful in principle, but is often out of date and so misleading. Tim Daly had been re-working Axiom to put much of its code in “literate” form [34] with the explicit objective of making it longer lasting: unfortunately this failed to lead to a big influx of new collaborators—it illustrates some of the challenges we are commenting on here.
@@ 3 Early Portability in Computer Algebra
Alpak was rewritten as a set of Fortran subroutines and paired with a language translator to become Altran[17] whose portability is mentioned by Johnson and Richie[32] when they were considering other software portability issues.
@@
The Camal[13] computer algebra system, developed by David Barton, Steve Bourne and John Fitch, was originally written in assembly code for the Titan computer at Cambridge. There were just two very similar computers in the world, and another three that were fairly close relatives.
Over the same period, George E. Collins was developing a set of libraries, PM [8], for big integer and polynomial arithmetic, first at IBM, then at the University of Wisconsin. We are under the impression that these began as assembly code libraries before the development shifted to Fortran, but would appreciate confirmation of this.
@@
Two alternative approaches emerged. The first was to rely on Lisp[47] as an implementation language, as Lisp had the concept of dynamic data structures at its core. The second was to rely on lower-level “systems programming languages” to provide efficient resource management. At the start this meant BCPL[43], although in due course C[33] claimed the dominant place in that “ecological” niche.
From its beginnings at the end of the 1950s, Lisp had matured and taken its place as a recognised family of closely related languages. There was no notion of a single, standard Lisp language. Important variants around 1970 were MacLisp[42], from MIT, and InterLisp[48], then from Bolt, Beranek and Newman. For some time, both of those had a clear enough common core that migrating software between them was reasonably sane for quite a while.
However by the end of the 1960s dialect divergence was sufficient to cause trouble and, Anthony Hearn introduced Standard Lisp[18] to serve as the foundation for the Reduce system. This was not so much intended for adoption outside Reduce usage, but it defined the expectations that Reduce had. It was consciously conservative and a bit minimal, and hence it was not especially attractive to those who wanted to develop and extend Lisp for direct use. So meanwhile the Macsyma system was built on MacLisp, first led by William Martin and then Joel Moses from 1968 to 1982 at MIT, and the Lisp dialect used there collected many many add-ons, where Joel Moses characterised the language as “a ball of mud”[39] that one could add arbitrary extras to it without altering its already slightly odd outline. Later, a port of Macsyma to the VAX architecture was achieved at UC Berkeley by creating a MacLisp-like base, Franz Lisp, by Richard Fateman and his students.
The Scratchpad system was built on Lisp/VM[3], led by Jim Griesmer at IBM. The IBM Lisp system was led by Fred Blair and later Cyril Alberga. (Note, this Scratchpad was retrospectively named Scratchpad I to distinguish it from the later, distinct system Scratchpad II which became Axiom.) Scratchpad (I) was seriously notable from the portability perspective because it started life by amalgamating major components from almost all of the Lisp-coded algebra projects extant around 1970, when the project was in its early stages and the various Lisp dialects were still close enough that code for one could be adapted for use on another with only a reasonable amount of effort.
@@
The different Lisp systems were not equally available on all platforms, so some systems made an attempt to run on more than one Lisp, despite the differences. Reduce is particularly notable in this regard. A version of it dated 1973 has been recovered and it can still be run — and amazingly much of the code in it can be recognised in the most recent versions of the system. This historical relic is available for download on the main Reduce website at https://sourceforge.net/projects/reduce-algebra. The strategy there was to rely on a minimal set of core Lisp functionality, and to reproduce necessary higher-level function as a set of Lisp functions relying only on that core. A similar approach was later adopted for Axiom.